OpenStudy (kylav210):

If the dimensions of a rectangle become one-fourth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle. Enter your answer in numerical form, using the / for the fraction bar if needed.

OpenStudy (anonymous):

Let the dimension be x and y.. So: \[Area = xy\] now both the dimensions have become 1/4 so: (1/4)x and (1/4)y \[Area(New) = \frac{1}{16}xy\] \[Area(New) = \frac{1}{16} (Previous \quad Area)\]

OpenStudy (kylav210):

Ohh okay! But if the whole rectangle is 1/4 smaller wouldn't the area be 1/4 smaller??

OpenStudy (kylav210):

@waterineyes

OpenStudy (anonymous):

Yeah it is smaller.. Suppose if original area is 80 then: \[Area(New) =\frac{1}{16} \times 80 = 5\]

OpenStudy (kylav210):

Got it ! Thank you!(:

OpenStudy (anonymous):

Welcome dear..